OpenStudy (anonymous):

Can someone please help me figure this out. It's die tomorrow.Add each of the following sets of vectors together using the component method. F1 = 140 N [W], F2 = 200 N [E30°N], F3 = 100 N [S35°W]

terenzreignz (terenzreignz):

First we need a convention... let's set north and east to be positive directions while setting west and south to be negative, ok?

OpenStudy (anonymous):

ok

terenzreignz (terenzreignz):

It's all a matter of adding components... Let's start with the first, shall we? \[\Large \left.\begin{matrix}FORCE & Horizontal & Vertical\\ F_1 & \color{red}?&\color{red}? \\ F_2 & & \\ F_3 & & \\ \color{red}{total} \end{matrix}\right.\]

terenzreignz (terenzreignz):

Now, if it was 140 newtons northwest, how much is its horizontal (in this case, westward) component?

terenzreignz (terenzreignz):



OpenStudy (anonymous):

oh okay. So I just have to find the components then add them up.

terenzreignz (terenzreignz):

Yup.

OpenStudy (anonymous):

Would the horizontal component would be 0? And vertical component be 140sin30

terenzreignz (terenzreignz):

For which? oh wait... F1 is just west, right? sorry... it's the vertical component that has to be zero if F1 is only going west -_-

OpenStudy (anonymous):

F1 is just West yup

OpenStudy (anonymous):

oh okay. sorry, physics is not my strong subject

terenzreignz (terenzreignz):

okay, let's key that in \[\Large \left.\begin{matrix}FORCE & Horizontal & Vertical\\ F_1 & \color{red}{140}&\color{red}0 \\ F_2 & & \\ F_3 & & \\ \color{red}{total} \end{matrix}\right.\]

terenzreignz (terenzreignz):

Is that right?

OpenStudy (anonymous):

Yeah. I do believe it would be

terenzreignz (terenzreignz):

LOL No it isn't :P Be very careful with signs... remember that west is negative (as is south) \[\Large \left.\begin{matrix}FORCE & Horizontal & Vertical\\ F_1 & \color{blue}{-140}&\color{blue}0 \\ F_2 &\color{red}? &\color{red}? \\ F_3 & & \\ \color{red}{total} \end{matrix}\right.\] Now what about F2?

OpenStudy (anonymous):

okay so F2 would equal to horizontal 200cos30=173.2 and vertical 200sin30=100

terenzreignz (terenzreignz):

Okay, yes :D \[\Large \left.\begin{matrix}FORCE & Horizontal & Vertical\\ F_1 & \color{blue}{-140}&\color{blue}0 \\ F_2 &\color{blue}{173.2} &\color{blue}{100} \\ F_3 &\color{red}? &\color{red}? \\ \color{red}{total} \end{matrix}\right.\]

OpenStudy (anonymous):

Is that right tho?

terenzreignz (terenzreignz):

Yup :)

OpenStudy (anonymous):

oh good okay. So F3 would have horizontal 100cos30=86.60 and vertical 100sin30=50

terenzreignz (terenzreignz):

Not right :D

terenzreignz (terenzreignz):

for one thing, it's 35 degrees in F3

OpenStudy (anonymous):

oops :p

OpenStudy (anonymous):

oh aahh

OpenStudy (anonymous):

okay so would it be horizontal 100cos35=81.9 and vertical 100sin35=57.36

OpenStudy (anonymous):

or do I have cosine and sine mixed up :p

terenzreignz (terenzreignz):

the horizontal doesn't always get the cosine... the one to the left of the angle measure (in this case, S) It's S35E So the one to the left is south, a vertical component... So, vertical gets cosine, horizontal gets sine. try again :P

terenzreignz (terenzreignz):

whoops, I meant S35W sorry

OpenStudy (anonymous):

Oh now I get it. :P Okay so it would be vertical 100cos35=81.9 and horizontal would be 100sine35=57.35. Is that right?

terenzreignz (terenzreignz):

Almost, but not quite. Do remember that these are south and west that we're dealing with...

OpenStudy (anonymous):

oh yes so it would have to be negative

OpenStudy (anonymous):

so it would be vertical -100cos35=-81.9 and horizontal would be -00sine35=-57.35.

terenzreignz (terenzreignz):

that's better :P

OpenStudy (anonymous):

-100sin35=-57.35

terenzreignz (terenzreignz):

\[\Large \left.\begin{matrix}FORCE & Horizontal & Vertical\\ F_1 & \color{blue}{-140}&\color{blue}0 \\ F_2 &\color{blue}{173.2} &\color{blue}{100} \\ F_3 &\color{blue}{-57.35} &\color{blue}{-81.9} \\ \color{red}{total} \end{matrix}\right.\]

terenzreignz (terenzreignz):

Now just add them up.

OpenStudy (anonymous):

alright so horizontal = to -24.15 and vertical adds up to 18.1.

terenzreignz (terenzreignz):

Is that all you need?

OpenStudy (anonymous):

I got it from here. Thank you so much. :)

terenzreignz (terenzreignz):

No problem :)

terenzreignz (terenzreignz):

Be very careful with signs... VERY CAREFUL LOL signing off --------------------------------- Terence out